OpenStudy (anonymous):

What is the simplest form of the number? -8^2/3 A. 4 B. 19 C. -9 D. -5.33

OpenStudy (anonymous):

\[-8^{\frac{ 2 }{3 }}=\sqrt[3]{-8^2}\]

OpenStudy (anonymous):

Ok o.o how does that help me find my answer?

OpenStudy (anonymous):

\[\sqrt[3]{-8^2}=\sqrt[3]{64}\] better?

OpenStudy (anonymous):

yes